OpenStudy (anonymous):

Will award a medal @ganeshie8

OpenStudy (anonymous):

i got 17/2

OpenStudy (anonymous):

my new transformation integral was \[1/2\int\limits_{0}^{3}\int\limits_{0}^{2}(-2/3u+7/3v)dudv\]

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (michele_laino):

why 1/2 ?

OpenStudy (michele_laino):

the Jacobian is 1/3

OpenStudy (michele_laino):

I got 17/3, is it right?

OpenStudy (anonymous):

i got the jacobian as 1/2 but that could be wrong.

ganeshie8 (ganeshie8):

with all the denominators as 3, there is no way for Jacobian to be 1/2

OpenStudy (anonymous):

1/6+2/6

ganeshie8 (ganeshie8):

you should be working 1/9 + 2/9

OpenStudy (anonymous):

ohghad 3x3=9

OpenStudy (michele_laino):

here is my Jacobian: \[\Large J = \left( {\begin{array}{*{20}{c}} {1/3}&{1/3} \\ { - 2/3}&{1/3} \end{array}} \right),\quad \left| J \right| = \left( {\frac{1}{3} \times \frac{1}{3}} \right) - \left\{ {\frac{1}{3} \times \left( { - \frac{2}{3}} \right)} \right\}\]

OpenStudy (michele_laino):

\[\Large \begin{gathered} J = \left( {\begin{array}{*{20}{c}} {1/3}&{1/3} \\ { - 2/3}&{1/3} \end{array}} \right),\quad \hfill \\ \hfill \\ \left| J \right| = \left( {\frac{1}{3} \times \frac{1}{3}} \right) - \left\{ {\frac{1}{3} \times \left( { - \frac{2}{3}} \right)} \right\} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

for this one, is the parameterization x=0,y=y,z=f(y)sin(theta)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (michele_laino):

hint: I rewrite your equation, as below: \[\large y = \frac{{{z^2}}}{{25}} + 1\] so we have: